OpenStudy (hockeychick23):

True or false. If the probability of success during a 4th event of a geometric experiment is 0.0729, the probability of success during the 5th experiment could be 0.06561. (I think it's true I just wanted to check, thanks!)

OpenStudy (tkhunny):

p(1) = p p(2) = qp p(3) = q^2*p p(4) = q^3*p = 0.0729 p(5) = q^4*p = 0.06561 Why do you think it is true?

OpenStudy (hockeychick23):

I think its true because^3*(1-p) = 0.0729 p^4*(1-p) = 0.06561 p = 0.06561/0.0729 = 0.9 and sinces 0≤p≤1, it is a valid probability

OpenStudy (hockeychick23):

and since 0<=p<=1 it is a valid probability

OpenStudy (hockeychick23):

or at least i think so